SOEST Home Search Site Map Links Contact
Dolphins at the bow of the Kilo Moana University of Hawaii Marine Center UH Marine Center Logo
Home Ships Marine Center Ship Techs Schedules Cruise Planning    

 

Overview

Layout

Handbook

Scientific Equipment

Support Equipment

Daily Report

Current Position

Schedule

Photo Album

 

Daily Report for R/V Kilo Moana May and June 2007


Greetings All,

KM-0708-07 Noon Report 1200 Saturday 30 June 2007

1200 Position: 22o-30.0 N 158o-00.0 W

Fuel on board:   68,278 gal.

Fuel consumed:  1,010 gal.

Wind is ESE at 22 kts with a moderate sea. The swell is E’ly at 6-7 ft. The Baro is steady at 1019 MLB.

The sky is partly cloudy with 3/10’s cover. The air temp is 81o F.

Station work was completed at Aloha Station at including an additional MET station at 0700 this morning. We are going to occupy three additional Met stations toward the South today and tonight. ETD last Met station at 0300 01 July. ETA Buoy “H” is 0900 and arrival Snug 0930. Our ETA is contingent on traffic moving in the harbor. As our original ETA for “H” Buoy was 1000. However, due to the necessity to off load WHOTS-4 Cruise upon arrival, it is desired to arrive at the buoy at 0900. The Chief Scientist is in complete agreement with our desire to arrive 1 hour earlier.

Warmest Regards,

Phil


Greetings All,

KM-0708-06 Noon Report !200 Friday 29 June 2007

1200 Position 220-45.0 N 158o-00.1 W

Fuel on board:    69,288 gal.

Fuel consumed:   1,058 gal.

Wind is ENE at 22 kts with a moderate sea. The swell is E’ly at 6-7 ft. The Baro is steady at 1019 MLB.

The sky is partly cloudy with 3/10’s cover. The air temp is 81o F.

We are at the center of Aloha Station conducting CTD casts. We had a deep cast this morning to 4700 meters. During the haul-in phase the wire was slushed by Chief mate and Dave Spurgin. All science operations have been going flawlessly. All engineering systems are performing to the Chief’s satisfaction. The food and preparation has been exceptional. In short, it has been a great cruise for me to be a part of.

Warmest Regards,

Phil


Greetings All,

KM-0708-05 Noon Report Thursday 28 June 2007

1200 Position 22o-43.6 N 158o-00.5 W

Fuel on board:   70,346 gal.

Fuel consumed:  1,150 gal.

Wind is ENE at 28 kts with a moderate sea. The swell is E’ly at 6-7 ft. The Baro is steady at 1021 MLB.

The sky is cloudy with 6/10’s cover. The air temp is 79o F.

The recovery of WHOT # 3 was completed at 1255 when the last of the hard hat floats were recovered. The total time for recovery was 6 hours and 10 minutes. All propulsion systems are working with no other problems noted. We are proceeding outside of Aloha Station to clear holding tank and refuse. CTD station work will resume at 0200 tomorrow morning at the center of Aloha Station. Today and thru the night we will occupy a position close aboard WHOT # 4 buoy for meteorological data comparison.

Warmest Regards,

Phil

 


Greetings All,

KM-0708-04 Noon Report 1200 Wednesday 27 June 2007

1200 Position 22o-45.4 N 157o-55.8 W

Fuel on board:   71,496 gal.

Fuel consumed:  1,050 gal.

Wind is ENE at 20 kts with a moderate sea, The swell is E’ly at 6-7 ft. The Baro is steady at 1022 MLB.
The sky is partly cloudy with 4/10’s cover. The air temp is 81o F.

Rich conducted ISM safety training for the crew in the use and practical demonstration of the “SKED” personnel recovery and evacuation devise. We are engaged in continuing CTD casts every four hours in the close proximity of WHOTS mooring number 3. The bow thruster problem seems to have been resolved, however about 30 minutes ago the stbd. main motor went down. At the time we were operating in Auto Head Joystick with the port and stbd. propulsion on line. The bow thruster was being used on occasion in manual control. The entire propulsion system was not being taxed in anyway when the stbd. motor went down. The Chief Engineer, Paul and Debby are looking into this particular problem. I’ll keep you advised on this particular problem…….

Warmest Regards,

Phil


Greetings all,

KM-0708-03 Noon report 1200 Tuesday 26 June 2007
1200 Position: 22o-40.2 N 155o-59.0 W

Fuel on board:   72,546 gal.
Fuel consumed:  1,061 gal.

Wind is ENE at 18 kts. with a moderate sea. The swell is E’ly at 6-7 ft. The Baro is steady at 1018 MLB.
The sky is partly cloudy with 4/10’s cover and the air temp is 77o F.

We are engaged in continuing CTD casts every four hours in the close proximity to WHOT 4 mooring. Science is going well. The engineers are continuing their trouble shooting of the continuing problem of the bow thruster tripping off line. It ran for a 30 minute period this morning under load during station keeping. But in the end, it tripped off line as before.

Warmest Regards,
Phil


 

Greetings All,

KM-0708-02 Noon Report 1200 Monday 25 June 2007
1200 Position: 22o-40.2 N 157o-58.7 W

Fuel on board:   73,607 gal.
Fuel Consumed:   1853 gal.

Wind is ENE at 15 kts with a slight sea. The swell is E’ly at 5-6 ft. The Baro is steady at 1017 MLB
The sky is cloudy with 8/10’s cover and the air temp is 77o F.

Vessel is towing the WHOT # 4 mooring array toward the drop site, 1.24 nautical miles distant. The deployment commenced at 0650.
We should be in position to set the anchor at around 1320. The engineers are still trouble shooting the bow thruster problem.
It continues to trip off line as before. Other than that, and the concerns of # 3 MPDG, the trip has been unremarkable….so far.

All science operations have been flawless.

Warmest Regards,
Phil

KM 0707 02 Noon Report 1200(W) Fri Jun 08 2007 
Noon Position is: 21-20.6 N 158-16.4

Fuel on board: per departure report
Fuel Consumed (24h):
Potable water on board: per departure report

WX Wind 125 @ 14 kts, Swell 2-4 ft from 125
Barometer 1018.2 falling, Temp 82F, RH 72%, scattered clouds  

Vsl is on Kaena Station conducting test cast and will then proceed to Aloha Station. 

Preventive maintenance pretty much completed on 01-level.  Preparing to address deck on 02-level upcoming import period.   Continued miscellaneous interior and exterior work including repairs to tile/underlayment, overhead panel cleaning and various hull trunk field days efforts.

Rick

 

Following are the particulars of KM 0707 arrival at Honolulu Mon Jun 11:

 

0830 Completed all pre-arrival gear test and inspection

0854 Passing “H” bouy inbound

0930 1st line ashore

0936  FWE/SM

 

Fuel on board: 57,853 gal

Potable water:  7300 gal

 

Following statistics reported for KM 0707:

 

Total U/W time: 3d 00h 42m 

Total miles:  330.2    

Fuel consumed:  3430 gal

 

Rick

 

KM 0707 04 Noon Report 1200(W) Sun Jun 10 2007 

Noon Position is: 22-47.2N 158-07.3W

 

Fuel on board: 57,853 gal

Fuel Consumed (24h): 1271 gal

Potable water on board: 8068 gal

 

WX Wind 070 @ 20 kts, Swell 4-6 ft from 080

Barometer 1018.0 steady, Temp 78F, RH 69%, overcast  

 

Vsl is at Aloha Station conducting modified operations of the HOT-192 cruise plan as a result of CTD winch failure which has compromised cast timeline.  Failure of the CTD is attributed to a frozen brake assembly compounded by possible electrical problem to control joystick.  Science has requested early arrival at Honolulu.

 

ETA “H” buoy 0900 Jun 11

 

Rick
KM 0707 03 Noon Report 1200(W) Sat Jun 09 2007 

Noon Position is: 22-45.0N 158-00.0W

 

Fuel on board: 59,516 gal

Fuel Consumed (24h): 1767 gal

Potable water on board: 7899 gal

 

WX Wind 095 @ 20 kts, Swell 4-6 ft from 090

Barometer 1018.0 falling, Temp 81F, RH 78%, scattered clouds  

 

Vsl is at Aloha Station conducting operations in accordance with HOT-192 cruise plan.  Science going well.

Rick

Following are the particulars of KM 0706 arrival at Honolulu Wed Jun 06:

0630 Completed all pre-arrival gear test and inspection
0657 Passing “H” bouy inbound
0741 1st line ashore
0800  FWE/SM

Fuel on board: 61,283gal
Potable water:  6124 gal

Following statistics reported for KM 0706:

Total U/W time: 26d 00h 18m 
Total miles:  5686.5    
Fuel consumed:  67,070 gal

Rick

KM 0706 28 Noon Report 1200(W) Mon Jun 05 2007 
Noon Position is: 20-53.2N 161-19.4W

Fuel on board: 63,561 gal
Fuel Consumed (24h):  2805 gal
Potable water on board:  6763 gal

WX Wind 070 @ 14 kts, Swell 4-6 from 060
Barometer 1018.3 stdy, Temp 81F, RH 82%, scattered clouds  

Vsl is underway conducting PLUME mapping enroute Hononlulu.  28 OBS units on board with 10 failing to communicate and or surface.  Science ready to go home.  Ship doing well. 

ETA Hotel Buoy 0700 06 June   

Rick

 

KM 0706 27 Noon Report 1200(W) Mon Jun 04 2007 
Noon Position is: 19-59.3N 162-01.6W

Fuel on board: 66,366 gal
Fuel Consumed (24h): 2822 gal
Potable water on board: 7380 gal

WX Wind 090 @ 13 kts, Swell 3-4 from 080
Barometer 1017.7 falling, Temp 83F, RH 75%, scattered clouds  

Vsl is on station waiting for last OBS to surface. 27 units on board. Science doing well, ship doing well. 

ETA Hotel Buoy 0700 06 June   

Rick

KM 0706 25 Noon Report 1200(W) Sat Jun 02 2007 
Noon Position is: 18-44.5N 159-07.5W

Fuel on board: 71,917 gal
Fuel Consumed (24h): 2836 gal
Potable water on board: 7522 gal

WX Wind 100 @ 10 kts, Swell 2-4 ft from 145
Barometer 1015.8 steady, Temp 82F, RH 74%, scattered clouds  

Vsl enroute to next station – ETA 0030.  25 units on board.  Science doing well, ship doing well. 

ETA Hotel Buoy 0700 06 June   

Rick

KM 0706 26 Noon Report 1200(W) Sun Jun 03 2007 
Noon Position is: 19-48.9N 160-03.3W

Fuel on board: 69,188 gal
Fuel Consumed (24h): 2729 gal
Potable water on board: 7980 gal

WX Wind 080 @ 11 kts, Swell - slight
Barometer 1016.8 steady, Temp 85F, RH 76%, scattered clouds  

Vsl is on station waiting for OBS to surface.  26 units on board.  One station remaining.  Science doing well, ship doing well. 

ETA Hotel Buoy 0700 06 June   

Rick

 

KM 0706 24 Noon Report 1200(W) Fri Jun 01 2007 
Noon Position is: 17-55.5N 160-21.5W

Fuel on board: 74,753 gal
Fuel Consumed (24h): 2162 gal
Potable water on board: 7170 gal

WX Wind 070 @ 12 kts, Swell 3-5 ft from 080
Barometer 1015.2 steady, Temp 82F, RH 72%, scattered clouds  

Vsl enroute to next station for daylight recovery tomorrow morning due to questionable beacon/strobe status of the unit.    24 units on board with 4 stations remaining.  Science doing well, ship doing well. 

Preventive maintenance continues on 01-level and other areas in need about the ship.  Ready to roll out aft portions of 01 deck port and stbd.   Using grinders, sanding disks and scrappers/chippers in lieu of needle guns for bulk of deck work.  Continued miscellaneous interior and exterior work including repairs to buckled tile/underlayment and overhead panel cleaning.

ETA Hotel Buoy 0700 06 June   

Rick

KM 0706 23 Noon Report 1200(W) Wed May 31 2007
Noon Position is: 16-50.0N 159-40.1W

Fuel on board: 76,915 gal
Fuel Consumed (24h): 2382 gal
Potable water on board: 7038 gal

WX Wind 080 17kts, Swell 3-5 ft from 080
Barometer 1015.6 falling, Temp 80F, RH 73%, scattered clouds  

Vessel is on station conducting cast with rescue beacon.  Communications was lost with this unit on deployment last year and extended station time had been scheduled for this recovery.  23 units on board.  Science doing well, ship doing well.        

Rick

KM 0706 22 Noon Report 1200(W) Wed May 30 2007
Noon Position is: 14-40.0N 157-52.1W

Fuel on board: 79,297 gal
Fuel Consumed (24h): 2185 gal
Potable water on board: 7431 gal

WX Wind 080 17kts, Swell 3-5 ft from 090
Barometer 1015.0 falling, Temp 80F, RH 70%, scattered clouds  

ETA next station 1900.   Science doing well with 22 units on board, 7 stations to go and 8 no-shows.  Ship doing well with intermittent bow thruster problems. 

Rick

KM 0706 20 Noon Report 1200(W) Tue May 29 2007
Noon Position is: 16-46.1N 157-39.5W

Fuel on board: 81,482 gal
Fuel Consumed (24h): 2689 gal
Potable water on board:  7736 gal

WX Wind 075 13kts, Swell 3-5 ft from 050
Barometer 1015.6 steady, Temp 80F, RH 72%, scattered clouds  

ETA next station 1400.   Science doing well.  Ship doing well. 

Rick

 KM 0706 19 Noon Report 1200(W) Mon May 28 2007

Noon Position is: 15-59.0N 155-52.0W

 

Fuel on board: 84,171 gal

Fuel Consumed (24h): 2725 gal

Potable water on board: 8000 gal

 

WX Wind 080 18kts, Swell 4-6 ft from 060

Barometer 1016.0 steady, Temp 79F, RH 60%, scattered clouds   

ETA next station 1430.   More OBS units finding their way onto the aft, bringing smiles to science.  Vessel doing well. 

 

Rick

KM 0706 18 Noon Report 1200(W) Sun May 27 2007 

Noon Position is: 14-26.5N 152-13.5W

 

Fuel on board: 86,896 gal

Fuel Consumed (24h): 2284 gal

Potable water on board: 7993 gal

 

WX Wind 070 @ 17 kts, Swell 5-8 ft from 080

Barometer 1014.8 steady, Temp 80F, RH 58%, scattered clouds  

 

ETA next station 2200 this evening.   Vessel doing well, science smiling after morning recovery. 

 

Rick

KM 0706 17 Noon Report 1200(W) May 26 2007 

Noon Position is: 15-57.0N 149-40.3W

 

Fuel on board: 89,180 gal

Fuel Consumed (24h): 2510 gal

Potable water on board: 7517 gal

 

WX Wind 070 @ 19 kts, Swell 6-9 ft from 070

Barometer 1015.0 falling, Temp 79F, RH 70%, scattered clouds  

 

ETA on station 1315.  Vessel doing well, science not doing so well.

 

Rick

 

KM 0706 16 Noon Report 1200(W) May 25 2007 

Noon Position is: 19-06.7N 149-18.2W

 

Fuel on board: 93,009 gal

Fuel Consumed (24h): 2150 gal

Potable water on board: 7730 gal

 

WX Wind 080 @ 18 kts, Swell 4-6 ft from 070

Barometer 1019.2 steady, Temp 78F, RH 71%, scattered clouds  

 

Vessel on station with ETD 1300.  Have experienced failed OBS returns at evening station of 24th and morning station of 25th.  First rate communications established on the evening station with deployed unit but the unit failed to release after multiple burn commands.  No communications has been established with the current unit. Science party understandably unhappy. 

 

Preventive maintenance continues on 01-level.  Spot work aft nearly completed in vicinity of CTD winch, NA crane and van areas.  Miscellaneous interior and exterior work ongoing with the ship beginning to “clean up” a bit.   

 

Rick

KM 0706 15 Noon Report 1200(W) May 24 2007 
Noon Position is: 18-03.5N 151-05.2W

Fuel on board: 95,159 gal
Fuel Consumed (24h): 2899 gal
Potable water on board: 7993 gal

WX Wind 110 @ 16 kts, Swell 5-8 from 080
Barometer 1018.7 rising, Temp 78F, RH 85%, scattered clouds  

Continue with recovery operations – ETA next station 1630.  Science doing well.  Ship doing well.

Rick 

 

KM 0706 14 Noon Report 1200(W) May 23 2007 
Noon Position is: 17-13.9N  153-46.1W

Fuel on board: 98,058 gal
Fuel Consumed (24h): 2484 gal
Potable water on board: 7349 gal

WX Wind 070 @ 19kts, Swell 6-9 from 110
Barometer 1016.8 steady, Temp 79F, RH 73%, scattered clouds  

Hump Day on the downhill side of retrievals.  15 OBS units on deck with 4 failing to communicate and surface.  Science doing well despite distress over failed instrument release.  Ship doing well.

Rick

KM 0706 12 Noon Report 1200(W) May 21 2007 
Noon Position is: 19-30.9N  153-58.0W

Fuel on board: 100,542
Fuel Consumed (24h): 1932
Potable water on board: 6239l

WX Wind 100 @ 17kts, Swell 4-6 from 120
Barometer 1017 falling, Temp 79F, RH 69%, scattered clouds  

ETA next station 1940.  Continue to have problems with communicating to deployed units.  Have pretty much removed ship noise as the culprit.  Science systems all tested and retested.  Ship is doing well.

Rick 

KM 0706 12 Noon Report 1200(W) May 21 2007 
Noon Position is: 20-36.4N 152-10.0W

Fuel on board: 102,475 gal
Fuel Consumed (24h): 2445 gal
Potable water on board: 6621 gal

WX Wind 100 @ 19 kts, Swell 3-5ft from 120
Barometer 1020 steady, Temp 79F, RH 72%, scattered clouds  

Anticipate arrival at next station 1210.  Unit failed to communicate and surface at yesterday evenings station.  Vessel will arrive next station 1210.  Science doing well, ship doing well.

Rick 

KM 0706 11 Noon Report 1200(W) May 20 2007 
Noon Position is: 22-23.7N 150-38.6W

Fuel on board: 104,919 gal
Fuel Consumed (24h): 2688 gal
Potable water on board: 6449 gal

WX Wind 130 @ 5 kts, Swell 3-5ft from 115
Barometer 1020 falling, Temp 75F, RH 71%, scattered clouds  

Vessel enroute to next station - ETA 2010.  Science doing well, ship doing well.

Rick 

KM 0706 10 Noon Report 1200(W) May 19 2007 
Noon Position is: 22-21.0N 152-43.8W

Fuel on board: 107,607 gal
Fuel Consumed (24h): 2527 gal
Potable water on board: 5700 gal

WX Wind 130 @ 17 kts, Swell 3-5ft from 150
Barometer 1020 falling, Temp 83F, RH 73%, scattered clouds  

Vessel completed recovery 1143 and is enroute to next station - ETA 2020.  Xcience doing well, ship doing well.

Rick 

KM 0706 09 Noon Report 1200(W) May 18 2007 
Noon Position is: 21-38.9N 154-57.0W

Fuel on board: 110,134 gal
Fuel Consumed (24h): 2602 gal
Potable water on board: 6208 gal

WX Wind 100 @ 14 kts, Swell 3-4 ft from 115
Barometer 1018.3 falling, Temp 77F, RH 77%, scattered clouds  

Vessel enroute to next station - ETA 1630.  Expecting recovery at or about 1900.  Science doing well, ship doing well.

Preventive maintenance continues on 01-level.  Deck force has been limited to 1 hour per day using needle guns but have redoubled our efforts and doing things the old fashion way with chipping hammers, scrappers and wire brush / grinders.  Completed spot work on areas previously inaccessible due to three embarked vans and addressing 01 port and stbd aft by CTD winch and NA crane. 

Rick

 

KM 0706 08 Noon Report 1200(W) May 17 2007 
Noon Position is: 22-42.7N 158-08.0W
Fuel on board: 112,736 gal
Fuel Consumed (24h): 2659 gal
Potable water on board: 6603 gal

WX Wind 100 @ 15 kts, Swell 3-4 ft from 110
Barometer 1017.6 falling, Temp 80F, RH 72%, scattered clouds  

Vessel approaching station – ETA 1230.  Expecting recovery at or about 1500.  Science doing well, ship doing well.

Rick

 

KM 0706 07 Noon Report 1200(W) May 16 2007 
Noon Position is: 24-12.1N 154-48.9W
Fuel on board: 115,395 gal
Fuel Consumed (24h): 2644 gal
Potable water on board: 6974 gal

WX Wind 090 @ 12 kts, Swell 3-4 ft from 160
Barometer 1020.1 steady, Temp 79F, RH 68%, scattered clouds  

Vessel on station waiting for OBS to surface – expected on surface at 1215.  ETA next station 2230 or thereabouts depending on recovery time.  Science doing well, ship doing well.

Rick

 

KM 0706 06 Noon Report 1200(W) May 15 2007 
Noon Position is: 26-02.5N 154-28.0W
Fuel on board: 118,039 gal
Fuel Consumed (24h): 2724 gal
Potable water on board: 7424 gal

WX Wind 100 @ 10 kts, Swell 3-4 ft from 320
Barometer 1021.4 falling, Temp 81F, RH 72%, scattered clouds  

Vessel enroute next OBS site – ETA 2000.  Recovered one unit early this morning.  Favorable weather for all recoveries has been most helpful.  Science doing well, ship doing well, cribbage tournament in play.

Rick

KM 0706 05 Noon Report 1200(W) May 14 2007 
Noon Position is: 26-32.4N 158-45.3W
Fuel on board: 120,763 gal
Fuel Consumed (24h): 1921 gal
Potable water on board: 7710 gal

WX Wind Light & Variable, Swell 2-3 ft from 050
Barometer 1018.5 steady, Temp 82F, RH 75%, scattered clouds  

Vessel enroute next OBS site – ETA 0140.  Two units failed to surface at last site.  One unit set last year from KM and the other from an earlier cruise.  Science doing well, ship doing well.

Rick

 

KM 0706 04 Noon Report 1200(W) May 13 2007 
Noon Position is: 24-48.5N 158-46.5W
Fuel on board: 122,684 gal
Fuel Consumed (24h): 2472 gal
Potable water on board: 7211 gal

WX Wind 115 @ 8 kts, Swell 3-5 ft from 040
Barometer 1017.8 steady, Temp 81F, RH 71%, scattered clouds  

Vessel enroute next OBS site where we will recover one and possibly a second unit - ETA 2200.  Science doing well, ship doing well.

Rick

KM 0706 03 Noon Report 1200(W) May 12 2007 
Noon Position is: 22-57.0N 161-12.4W
Fuel on board: 125,156 gal
Fuel Consumed (24h): 2533 gal
Potable water on board: 7226 gal

WX Wind 100 @ 14 kts, Swell 3-5 ft from 040
Barometer 1017.0 steady, Temp 75F, RH 70%, scattered clouds  

Vessel enroute to Sta 3 with an ETA of 1715.  Two successful recoveries in fair weather conditions averaging 19 min per recovery.  Stern control station playing out nicely.  Science appears happy and ship doing well.

Rick

KM 0706 02 Noon Report 1200(W) May 11 2007 
Noon Position is: 21-17.0N 158-39.6W
Fuel on board: Per departure report
Fuel Consumed (24h): N/A
Potable water on board: Per departure report

WX Wind 050 @ 9 kts, Swell 3-5 ft from 110
Barometer 1016.9 falling, Temp 79F, RH 70%, scattered clouds  

Vessel enroute to 1st array site.  ETA 1300.

No major deck maintenance accomplished due to science ops, fueling, crew changeout, discharge and loadout.  Hull cleaned of rust stains, minor touchup painting completed.   

Rick

ollowing are the particulars of KM 0705 arrival at Honolulu May 07:

0615 Completed all pre-arrival gear test and inspections
0639 Passing “H” bouy inbound
0724 1st line ashore
0732  FWE/SM @ Pier 30 for fuel

Fuel on board: 36200 gal
Potable water: 6400 gal

Following statistics reported for KM 0705:

Total U/W time: 3d 22h 24m 
Total miles:   387.4  
Fuel consumed: 5928 gal

Rick

 

KM 0705 05 Noon Report 1200(W) May 06 2007

Noon Position is: 22-41.2N 158-00.0W

Fuel on board: 37485 gal

Fuel Consumed (24h): 1452 gal

Potable water on board: 6878 gal

 

WX Wind 110 @ 14 kts, Swell 4-6 ft from 090

Barometer 1017.2 steady, Temp 79F, RH 72%, scattered clouds

 

Vessel is on Aloha Station conducting operations per HOTS-191 General Cruise
Plan.  Science going well.

 

Rick

KM 0705 04 Noon Report 1200(W) May 05 2007

Noon Position is: 22-46.8N 157-58.1W

Fuel on board: 38937 gal

Fuel Consumed (24h): 1232 gal

Potable water on board: 7288 gal

 

WX Wind 120 @ 14 kts, Swell 3-5 ft from 100

Barometer 1016.2 steady, Temp 79F, RH 71%, scattered clouds

 

Vessel is on Aloha Station conducting operations per HOTS-191 General Cruise
Plan.  Science going well.

 

Rick

KM 0705 03 Noon Report 1200(W) May 04 2007

Noon Position is: 22-44.8N 158-00.3W

Fuel on board: 40169 gal

Fuel Consumed (24h): 1959 gal

Potable water on board: 7401 gal

 

WX Wind 110 @ 17 kts, Swell 4-5 ft from 090

Barometer 1016.0 steady, Temp 65F, RH 66%, scattered clouds

 

Vessel is on Aloha Station conducting operations per HOTS-191 General Cruise
Plan.  Science going well.

 

No maintenance to report for past week due to arrival, offload, crewchange,
onload, departure.

Rick

 



Archives - March and April 2007

Archives - January and February 2007

Archives - December 2006

Archives - November 2006

 

 

 


School of Ocean and Earth Sciences and Technology, University of Hawaii

University of Hawaii Marine Center

1 Sand Island Access Road, Honolulu, HI 96819

Phone (808) 842-9813, Fax (808) 842-9833 Toll Free 1-888-800-0460

email snug@soest.hawaii.edu

University of Hawaii Link